Give the formula for class mark and class size of a data.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Hint: To get the formula for the class mark and class size, just follow their definitions. The class mark is the average of the upper and lower limit of a class interval and the class size is the difference of the upper and lower limit of a class interval. Complete step by step answer: According to the question, we have to get the formula for the class mark and class size. We know that class mark is the mid-value or central value of a class. It means that it is the average of the actual upper limit and actual lower limit for a given class. It may also be said that the mid value of every class interval is called its class mark. The class mark can be calculated by the formula, ClassMark=actualupperlimit+actual lowerlimit2 …………………………………………………..(1) For instance, let us consider the class interval 10-20, The actual upper limit of the given class = 20 ………………………………….(2) The actual lower limit of the given class = 10 ……………………………………(3) Now, from equation (1), equation (2), and equation (3), we get ClassMark=20+102=15 Therefore, the class mark for the class interval 10-20 is 15. We also know that the class size is defined as the difference between the actual upper limit and actual lower of a given class interval. The class size can be calculated by the formula, classsize=actualupperlimit - actuallowerlimit ……………………………………(4) For instance, again let us consider the class interval 10-20, The actual upper limit of the given class = 20 ………………………………….(5) The actual lower limit of the given class = 10 ……………………………………(6) Now, from equation (4), equation (5), and equation (6), we get Classsize=20−10=10 Therefore, the class size for the class interval 10-20 is 10. From equation (1) and equation (4), we have formula for the class mark and class size for a given class interval respectively. Hence, ClassMark=actualupperlimit+actual lowerlimit2 and classsize=actualupperlimit - actuallowerlimit . Note: In this question, one might get confused and consider class size as the summation of the upper and lower limits of a given class interval. This is wrong because class size is the difference between the upper and lower limits of a given class interval. Therefore, one must keep this point in mind.